OpenStudy (anonymous):

lim of cos^2(x)/x as x approaches -infinity

OpenStudy (anonymous):

Since \[cos^2(x) \] is definitely going to remain finite \[\frac{cos^2(x)}{x}\] it going to hit 0 because any finite number divided by infinity is goin to be 0. Now to actually show this, I would recommend something like squeeze theorem. A delta epsilon proof would be pretty tricky.

OpenStudy (anonymous):

so basically 0/x<=cos^2(x)/x<=1/x

OpenStudy (anonymous):

Good job! Actually I wasn't sure which functions to use, but those functions definitely should work!

OpenStudy (anonymous):

so would lim(x approches infinity) (sqrt(x^2+1))/(2x+1) be DNE since it would basically be infinity/infinity?